ADNAN MUHAMMED

Grade 12,

In triangle ABC r = 1 R=3 s=5 then area of triangle ABC

In triangle ABC r = 1 R=3 s=5 then area of triangle ABC

Grade:11

1 Answers

Arun
6 years ago
Dear Pranit
 
Area of triangle (Delta) = s *r 
Area = 5*1 = 5 unit²
 
Regards
Arun (askiitians forum expert)